OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (4, –2) and has a slope of –3? OPTIONS: A)x + 3y = 10 B) x + 3y = –10 C) 3x + y = 10 D) 3x + y = –10

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(4\quad ,&-2) \end{array} \\\quad \\ slope = m= -3 \\ \quad \\ y-y_1=m(x-x_1)\quad \textit{plug in your values and solve for "y"}\)

OpenStudy (anonymous):

i dont get how to solve foy y

OpenStudy (jdoe0001):

well, like you did on the previous one, you isolate "y"

OpenStudy (jdoe0001):

hmmm actually the choices show that you don' t have to :/

OpenStudy (anonymous):

y-(-2)=-3(x-4) y+2=-3x+12 y=3x+10

OpenStudy (jdoe0001):

y -(-2) = -3 ( x - 4) y+2 = -3x+12 3x + y = 10

OpenStudy (anonymous):

okay thanks

OpenStudy (jdoe0001):

pretty much, yes, yours is correct, just missed the - in the last line

OpenStudy (anonymous):

y-(-2)=-3(x-4) y+2=-3x+12 y=3x+10

OpenStudy (anonymous):

i dont get how to solve foy y